OpenStudy (anonymous):

solve 4^(2x)-20*4^x+64=0

OpenStudy (anonymous):

solve \[u^2-20u+64=0\] and then replace u by \[4^x\] if the solution is negative, ignore it because \[4^x>0\]

OpenStudy (anonymous):

wait, what about u?

OpenStudy (anonymous):

\[u^2-20u+64=0\] \[(u-16_)(u-4)=0\] \[u=16,u=4\] \[4^x=16\iff x=2\] \[4^x=2\iff x=\frac{1}{2}\]

OpenStudy (anonymous):

isnt the second one 4 not 2?